OpenStudy (anonymous):

Debra earns $7.50 an hour, plus tips, as a waitress. Suppose she works 24 hours in one week. How much money must she receive in tips to earn more than $250 that week?

OpenStudy (anonymous):

f(x) = 7.5(x) f(24) = 7.5 * 24 f(24) = 180 250 - 180 = tips

OpenStudy (anonymous):

im working with inequalities so would the < with line under it be correct

OpenStudy (anonymous):

No, just <.